# RaLiBEV: Radar and LiDAR BEV Fusion Learning for Anchor Box Free Object Detection Systems

### 01 Dec 2023 v4

@@ III-C Interactive Transformer for Fusion of Radar and LiDAR in BEV
In the pursuit of enhancing the performance of sensor fusion in autonomous driving systems, we propose a novel approach that transcends the conventional method of simply concatenating features from each sensor. The traditional approach, while straightforward, is inherently limited in its ability to effectively weigh the features from each sensor. To address this limitation, we have designed an innovative interactive transformer-based BEV Fusion block. This block leverages modifications to the standard transformer, as described in [51], to fuse data from the two sensor branches more effectively. Fig. 4 provides a visual representation of two distinct fusion logics that we have developed.
The first approach Direct Interactive Transformer for BEV Fusion, depicted in Fig. 4(a), employs a direct interactive BEV fusion. This method utilizes one sensor branch feature map as the query and the other as the key and value. The weight matrix is computed through matrix multiplication of the query and key, which is subsequently normalized using a softmax function. The final output is the dot product of the weight and value, where the value is the LiDAR feature map. This process can be mathematically represented as follows:
$\bm{X}_{out}=\sigma(\bm{X}_{s_{0}}\bm{X}_{s_{1}}^{T})\odot\bm{X}_{s_{1}}+\bm{X}_{s_{1}}.$ (15)
In this equation, $\bm{X}$ represents the feature map, $\sigma(\cdot)$ denotes the softmax function, and $s_{0}$, $s_{1}$ are different sensors serving the roles of query and key.
Our second approach shown in Fig. 4(b), Dense Query Map-based Interactive Transformer for BEV Fusion (DQMITBF), shares the core concept with the first approach but introduces a variation in the query definition. In this method, we define a randomly initialized learnable query that shares the same size as the input key feature map. The interactive attention is computed between this learnable query and the concatenated radar and LiDAR feature map, which is treated as the key. This process can be mathematically expressed as:
$\bm{X}_{out}=\sigma(\bm{X}_{Q}\bm{X}_{C}^{T})\odot{\bm{X}_{C}}+\bm{X}_{C}.$ (16)
In this equation, $\bm{X}_{Q}$ is the randomly initialized learnable query, and $\bm{X}_{C}$ is the concatenated feature map of radar and LiDAR.
These innovative approaches to sensor fusion offer a more effective way to weigh and combine features from different sensors, thereby enhancing the performance of autonomous driving systems.
@@ IV-A Dataset and Evaluation Metrics
The synchronization issue between LiDAR data and radar data and the annotation generation was solved by interpolation with the help of ego vehicle pose [24]. Specifically, the pose parameters compensated for the misalignment between the LiDAR point cloud and radar range-azimuth heatmap. Utilizing the timestamps as a reference, [24] computes the motion between LiDAR point cloud and radar heatmap, then transfer every point to radar coordinate. The annotation bounding boxes are also created by [24] every 20 frames. The remaining frames’ annotations were also generated by interpolation. For testing in a foggy scenario, MVDNet relocates the LiDAR point cloud with the method of DEF [48]. According to the fog model in DEF, MVDNet sets the fog probability to 0.5 to recompute a maximum visible range for every point. If the current point range is greater than the maximum foggy range, this point would be set rather lost or relocated as a scatter point. Following the same operations on data in MVDNet, the 2D label bounding boxes are used as supervision. Corresponding radar and LiDAR data in 8862 number frames are split into 7071 for training and 1791 for testing.
@@ IV-C Experimental Results and Analysis
Direct Interactive Transformer for BEV Fusion 97.8 97.5 92.3
@@
TABLE III: Comparison of AP of oriented bounding boxes in bird’s eye view between our proposed RaLiBEV and other state-of-the-art radar and LiDAR fusion-based methods. The numbers in bold are the best of all methods.
@@
TABLE IV: Comparison of LiDAR only approaches and RaLiBEV which employs fusion of radar and LiDAR. The numbers in bold are the best of all methods.
Method Train Modality Clear+Foggy Clear-only
Test Clear Foggy Clear Foggy
IoU 0.5 0.65 0.8 0.5 0.65 0.8 0.5 0.65 0.8 0.5 0.65 0.8
Figure 5: Visualization of detection results of MVDNet and RaLiBEV.
@@ IV-C1 Ablation Study on Label Assignment Strategies
The second type of label assignment strategy is DIPS Assignment. This strategy presents improvement not much at the low precision AP result at IoU of 0.5 and 0.65, about 0.7% - 0.9%. Still, better performance at high precision AP at IoU of 0.8 results about 6% compares to the first strategy above. The result of the third strategy illustrates the effectiveness of choosing a more appropriate positive-sample strategy. Compared to the previous strategy, the performance of GAHPS Assignment shows another performance improvement of about 5% in high precision, which reaches 89.5% AP at IoU of 0.8. Following this idea, heatmap plus IoU cost design are implemented to upgrade the previous strategy. In GAHIPS Assignment, the performance at IoU of 0.5 maintains almost the same as the previous strategy, while the most notable change is on AP at IoU of 0.8. The performance beaks through to 92.0%, which is a great improvement of 10.5% compared to 81.5% of state-of-the-art ST-MVNDet++.
@@
The above single positive label assignment strategies are based on propelling the regression loss to approach the classification loss. These type of ideas shows great validity. In this strategy, the strong binding strategy has been experimented with. GACHIPS Assignment inherits the previous strategy, using heatmap and IoU as cost, calculating classification loss and regression loss after finding the best-matched anchor point. The final result shows further improvement at AP of IoU 0.5 and 0.8, which increased to 97.4% and 93.9%, respectively.
The multi-positive sample strategy designed in image object detection is actually to solve problems such as missed detection. The introduction of more positive sample points makes the model better distinguish the target and the background. However, the bounding boxes belonging to different targets do not overlap in the BEV perspective detection problem. Therefore, there is no problem that some targets have fewer pixels and are difficult to identify as foreground due to overlapping targets. Introducing multiple positive samples into BEV object detection does not consider the problems it brings to the regression, that is, the reduction of regression accuracy of the model. The experimental results show that the prediction regression results of many key points identified as positive samples have a small overlap with the ground-truth box. They are located far from the center point of the ground-truth box and have prediction boxes of the same size as the ground-truth box, but there is a larger offset. The performance of this point in statistics indicates that the low-precision AP results are not much different, while the high-precision AP@IOU=0.8 result is 18.2% lower than GACHIPS. This indicates that these positive sample points have a large deviation in the loss of $l_{\triangle xy}^{reg}$ and cannot continue to optimize. This is because the offset supervision values of multiple positive samples at different key point positions are variant. The farther away from the center point, the larger the offset supervision value is. These positive sample points of the same target provide the overall loss with different gradient values. However, the gradient descent algorithm can only regress each channel of each positive sample point to the same numerical level. Positive sample points with larger offset positions need higher weights to make their predicted values closer to the ground-truth value. However, to solve such problems introduced by the multi-positive sample design, the benefits are not as good as directly using the single-positive sample point design because focal loss can solve the problem of foreground-background classification in the current scene.
In the single positive sample design, a classification-guided regression strategy and classification-regression binding strategy are adopted to synchronize the contribution of classification and regression to the model’s gradient. Experiments show that the original DIPS strategy defines a set of optimal positive sample points for classification and regression through artificial methods. The classification and regression of positive samples are at the center of the supervision frame. This set of artificial positive points circumvents the sample assignment operation but suffers from limited performance. Based on classification-guided regression, the GAHPS and GAHIPS experiments show better performance. Based on the training classification results of the current round, the two strategies define the position of the regression positive sample points. As the training progresses, the regression of positive sample points of the model will gradually approach the positive sample points of classification. Unlike DIPS, because the model will regress key points other than the center point during the training process, the model has an ideal situation where the classification does not completely converge to the target center point. In other words, the classification score of the center point is not necessarily the highest. Therefore, after certain epochs of training, the model will infer the current best prediction bounding box based on the current results in the testing stage, which has the highest classification score and the smallest regression bias. In addition, it is found from the GAHIPS experimental results that when the regression prediction results are involved in sample assignment, the high-precision performance of the model is further improved. This indicates that the rotation IOU participates in model training guidance through sample assignment, further improving the model’s performance. Based on this, we believe that the center point assumption should be discarded, and the allocation search of positive and negative samples in training should be carried out in a completely free manner. GACHIPS takes the sum of the heatmap and IOU decoded by each candidate point in the Gaussian region as the cost and finds the best candidate point as the positive sample point of the current target to guide the model training and maximize the consistency of classification and regression. Experiments show that this method further improves the model’s overall performance and achieves a 1.9% improvement in high-precision performance.
@@ IV-C2 Ablation Study on Interactive Transformer-based BEV Fusion Approaches
To illustrate the effectiveness of interactive transformer-based BEV fusion approaches, the fusion module is inserted into the basic network framework with DIPS label assignment. Table II presents the results of transformers as well. In the Direct Interactive Transformer for BEV Fusion, the performance rapidly reached the same level as GACHIPS with only slightly lower performance, around 1.6% in high precision IoU of 0.8. The DQMITBF exhibits stronger performance. All three indicators exceed GACHIPS in an all-around way. This brings the AP result of IoU 0.5, 0.65, and 0.8 to 98.8%, 97.7%, and 94.8%, respectively.
To highlight the benefits of multi-modality fusion, Table IV presents a comparison of performance metrics between the LiDAR-only approach and the fusion of radar and LiDAR. It is evident from the comparison that neither RaLiBEV with LiDAR only nor PointPillars can match the performance levels achieved by the fusion of radar and LiDAR. This performance enhancement is particularly noticeable in foggy testing scenarios. When the model is trained on a “Clear+Foggy” dataset, RaLiBEV, in the radar and LiDAR fusion modality (denoted as “R+L”), exhibits exceptionally high performance, especially under the stringent requirements of a 0.8 IoU in “Foggy” test scenarios.
The results of this experiment show that a reasonable fusion logic would significantly enhance the performance of our RaLiBEV. Simple cross-attention between feature maps can increase performance. Still, query-based ideas can result in better effects depending on the flexibility of weights brought in by the learnable query.
@@ IV-C3 Performance Comparison with State-of-the-Art Methods on Different Weather Conditions
Following MVDNet [24], model performance in different weather is tested in this paper as described in Table III. The simulated foggy LiDAR data is generated by the code provided by MVDNet. This experiment is set as training using clear-only and clear plus foggy LiDAR and radar data, then testing under both clear and foggy data. Two peak performance models are chosen to prove their capability under this condition. One is RaLiBEV with GACHIPS, the other is RaLiBEV with DQMITBF. The performance of GACHIPS training under clear-only data and testing under the foggy scenario is better than DQMITBF’s. The performance of the horizontal comparison model after data enhancement is about 10%. GACHIPS increases from 84.6%, 82.4%, and 73.8% to 97.8%, 96.7%, and 93.7% under the foggy test. This result exceeds the ST-MVDNet [25] over 20% at clear plus foggy training and foggy test scenario. Compared to the newest work Bi-LRFusion [53], RaLiBEV outperforms a large margin of 20.7% in IoU 0.8 requirement. Meanwhile, according to the results of the state-of-art work ST-MVDNet++ [52], RaLiBEV has made progress of 4.4%, 7.8% and 19.0% in IoU 0.5, 0.65 and 0.8 under “Clear+Foggy” training, “Foggy” test. This result represents that RaLiBEV is able to perform better accuracy in adverse weather such as fog. The fusion logic and label assignment strategy help the detector achieves higher precision.
Fig. 5 shows the ground-truth, the estimation from MVDNet in clear weather, the estimation from RaLiBEV in clear weather, and the estimation from RaLiBEV in foggy weather, at the $3332^{th}$ and $4103^{th}$ frame of test dataset. The ground-truth bounding boxes are decorated with yellow color, and the predicted boxes by each model are shown with red color. All boxes use green lines to indicate the heading of the object. It is easy to find out that the estimations of MVDNet have 3 miss detection and 4 predicted boxes with reversed heading direction in the top row figure, and 1 false alarm and 2 reversed predicted boxes in the bottom row figure, though the weather is clear. In contrast, The RaLiBEV shows perfect detection results in clear weather, and only gives one false detection and no reversed heading estimation in both time frame under foggy weather.

### 06 Feb 2024 v5

@@ III-C Interactive Transformer for Fusion of Radar and LiDAR in BEV
In the pursuit of enhancing the performance of sensor fusion in autonomous driving systems, we propose a novel approach that transcends the conventional method of simply concatenating features from each sensor. The standard transformer [51] based traditional approach, while straightforward, is inherently limited in its ability to effectively weigh the features from each sensor. To this end, we have designed an innovative interactive transformer-based BEV Fusion block, to fuse data from the two sensor branches more effectively.
The first approach Direct Transformer for BEV Fusion, depicted in Fig. 4(a), employs a plain transformer BEV fusion as a basic method. This method utilizes radar branch feature map as the query and LiDAR feature map as the key and value. The weight matrix is computed through matrix multiplication of the query and key, which is subsequently normalized using a softmax function. Then the dot product of the weight and value are added with value itself acting as a residual process. This process can be mathematically represented as follows:
$\bm{X}_{out}=\sigma(\bm{X}_{R}\bm{X}_{L}^{T})\odot\bm{X}_{L}+\bm{X}_{L}.$ (15)
In this equation, $\bm{X}$ represents the feature map, $\sigma(\cdot)$ denotes the softmax function, $\odot$ denotes element-wise multiplication, and $\bm{X}_{R}$, $\bm{X}_{L}$ are radar and LiDAR feature maps. This method serves as a baseline to demonstrate the effectiveness of our second approach.
[…]
$\bm{X}_{out}=(\sigma(\bm{X}_{Q}\bm{X}_{R}^{T})\odot{\bm{X}_{L}}+\bm{X}_{L})\textcircled{c}(\sigma(\bm{X}_{Q}\bm{X}_{L}^{T})\odot{\bm{X}_{R}}+\bm{X}_{R}).$ (16)
The DQMITBF architecture interacts symmetrically with radar and LiDAR features through a BEV query and integrates them in a unified phase. This interactive query allows the model to prioritize feature-rich branches, especially in varying weather conditions. Thus, this novel approach optimizes the weighting and combination of heterogeneous sensor features in an adaptive manner, enhancing the effectiveness of autonomous driving systems in different meteorological scenarios.
@@ IV-A Dataset and Evaluation Metrics
The synchronization issue between LiDAR data and radar data and the annotation generation was solved by interpolation with the help of ego vehicle pose [24]. Specifically, the pose parameters compensated for the misalignment between the LiDAR point cloud and radar range-azimuth heatmap. Utilizing the timestamps as a reference, [24] computes the motion between LiDAR point cloud and radar heatmap, then transfer every point to radar coordinate. The annotation bounding boxes are also created by [24] every 20 frames. The remaining frames’ annotations were also generated by interpolation. For testing in a foggy scenario, MVDNet relocates the LiDAR point cloud with the method of DEF [48]. According to the fog model in DEF, MVDNet sets the fog probability to 0.05 to recompute a maximum visible range for every point. If the current point range is greater than the maximum foggy range, this point would be set rather lost or relocated as a scatter point. Following the same operations on data in MVDNet, the 2D label bounding boxes are used as supervision. Corresponding radar and LiDAR data in 8862 number frames are split into 7071 for training and 1791 for testing.
@@ IV-C Experimental Results and Analysis
Direct Transformer for BEV Fusion 97.8 97.5 92.3
@@
TABLE III: Comparison of LiDAR only approaches and RaLiBEV which employs fusion of radar and LiDAR. The numbers in bold are the best of all methods.
Method Train Modality Clear+Foggy Clear-only
Test Clear Foggy Clear Foggy
IoU 0.5 0.65 0.8 0.5 0.65 0.8 0.5 0.65 0.8 0.5 0.65 0.8
TABLE IV: Comparison of AP of oriented bounding boxes in bird’s eye view between our proposed RaLiBEV and other state-of-the-art radar and LiDAR fusion-based methods. The numbers in bold are the best of all methods.
@@
Figure 5: Interactive transformer feature map visualization results.
@@ IV-C1 Ablation Study on Label Assignment Strategies
The second type of label assignment strategy is DIPS Assignment. The improvement presented by this strategy is not significant at the low precision AP result at IoU of 0.5 and 0.65, approximately around 0.7% - 0.9%. Still, better performance at high precision AP at IoU of 0.8 results about 6% compares to the first strategy above. The result of the third strategy illustrates the effectiveness of choosing a more appropriate positive-sample strategy. Compared to the previous strategy, the performance of GAHPS shows another performance improvement of about 5% in high precision, which reaches 89.5% AP at IoU of 0.8. Following this approach, in GAHIPS, the performance at an IoU of 0.5 remains almost the same as the previous strategy, while the most notable change is in AP at an IoU of 0.8. The performance breaks through to 92.0%, which is a significant improvement of 10.5% compared to the 81.5% of the state-of-the-art ST-MVNDet++.
@@
The above single positive label assignment strategies are based on encouraging the best box regression predicted key point to approach the classification predicted key point. These types of ideas have shown great validity. However, in the following strategy, a strong binding strategy has been experimented with. GACHIPS inherits the previous strategy, using heatmap and IoU as cost, while calculating foreground-background classification loss and box loss after finding the best-matched anchor point. The final result shows further improvement at AP of IoU 0.5 and 0.8, which increased to 97.4% and 93.9%, respectively.
Experiments indicate that a multi-positive sample strategy can reduce a model’s high-precision performance. The reason is that calculating multiple losses for the same ground truth object introduces competition between samples. Specifically, the current loss design includes $\triangle{x}$ and $\triangle{y}$, which are related to the sample’s position. These supervision values vary by location, with points further from the center having larger values. This means that different positive samples for the same ground truth object have inconsistent regression endpoints, preventing the model from converging well near the extremes. In traditional image detection, multi-positive sample assignment strategies are popular for improving detection rates under overlap conditions. However, this approach is not suitable for BEV perception in autonomous driving, as targets do not overlap each other in BEV.
However, the assumption of a single positive sample point effectively avoids this competition, creating a singular convergence point for the model. Yet, inconsistencies between foreground-background classification and bounding box regression with a single positive sample point can limit the model’s ultimate performance. Specifically, when employing the DIPS strategy, the model uses a fixed target center point to calculate box regression loss, but the optimal point for foreground-background classification during inference may not be at the target center, leading to a loss in model accuracy. This design has a typical inconsistency issue between classification and regression. Therefore, the GAHPS strategy aims to find a way to ensure that the best foreground point during inference is also the one that has undergone the most box regression training by the end of training. Hence, during each training session, the peak points of the forward inference heatmap are selected as the anchor points for box regression. This strategy indeed brought an improvement in model performance. Compared to DIPS, its high-precision IoU 0.8 performance increased by 5%.
After observing the aforementioned performance improvements, we hypothesized that a better cost design would further enhance model accuracy. Thus, we replaced the original heatmap with a heatmap plus IoU, hoping to further reduce inconsistencies. Experiments show that this understanding is correct, and the model indeed achieved further improvements in high-precision performance. Compared to GAHPS, GAHIPS improved by 2.5% at IoU 0.8. Furthermore, we attempted to have the selected positive sample control both the foreground-background classification and the bounding box regression loss, proposing GACHIPS. This modification means that we have theoretically maximized the consistency between above two calculations. Specifically, during the training phase, the model uses the best anchor point obtained from inference for loss calculation. Then, during inference, it decodes the best detection box through this optimal anchor point. This design achieves unity between foreground-background classification and bounding box regression, ensuring consistency between the training and inference stages. Experiments show that this approach achieved the best high-precision performance, reaching 93.9% at IoU 0.8.
@@ IV-C2 Ablation Study on Interactive Transformer-based BEV Fusion Approaches
To illustrate the effectiveness of interactive transformer-based BEV fusion approaches, both the direct transformer and interactive transformer are applied in fusion module with DIPS label assignment, respectively. Table II presents the results of transformers as well. In the Direct Transformer for BEV Fusion, the performance reached the same level as GACHIPS with only slightly lower performance, around 1.6% in high precision IoU of 0.8. While the DQMITBF exhibits stronger performance, all three indicators exceed GACHIPS in an all-around way. This brings the AP result of IoU 0.5, 0.65, and 0.8 to 98.8%, 97.7%, and 94.8%, respectively.
To demonstrate the effectiveness of DQMITBF in a more comprehensible manner, the feature maps of two selected two frames in the interactive transformer fusion module are visualized as depicted in Fig. 5. In Fig. 5, The top two rows are corresponding to Frame 2799 and the bottom two rows are corresponding to Frame 3299, which are selected randomly. Each frame includes selections from clear and foggy weather conditions for analysis. For each column, the feature maps $\bm{X}_{Q},\bm{X}_{R},\bm{X}_{L},\bm{X}_{Q}\bm{X}_{R}^{T}$, and $\bm{X}_{Q}\bm{X}_{L}^{T}$ are illustrated for visual comprehension. Furthermore, all feature maps are normalized to the range $[0,1]$ for a consistent representation.
Utilizing a unified BEV query $\bm{X}_{Q}$, the input radar feature map $\bm{X}_{R}$ and LiDAR feature map $\bm{X}_{L}$ are weighted to $\bm{X}_{Q}\bm{X}_{R}^{T}$ and $\bm{X}_{Q}\bm{X}_{L}^{T}$ respectively, subsequently producing a fused feature for additional processing. It can be observed that $\bm{X}_{Q}$ manages radar and LiDAR features differently under varying weather conditions. In clear weather, $\bm{X}_{Q}\bm{X}_{L}^{T}$ is noticeably larger in numerical value than $\bm{X}_{Q}\bm{X}_{R}^{T}$. However, in foggy conditions, the input $\bm{X}_{L}$ is significantly impacted by noise, resulting in $\bm{X}_{Q}\bm{X}_{R}^{T}$ being larger than $\bm{X}_{Q}\bm{X}_{L}^{T}$. From these observations, we can draw the following conclusions:
  • The approach utilizes an interactive transformer to aggregate data from LiDAR and radar branches in a unified manner;
  • The BEV query in the interactive transformer can focus on more information-rich signals under different weather conditions.
Additionally, as depicted in Fig. 5, the input $\bm{X}_{L}$ and $\bm{X}_{R}$ appear visually similar to the raw source data, and $\bm{X}_{Q}$ does not seem to focus on a specific area determined by the presence of an object. This is because the interactive transformer fusion module operates in the encoder process of the entire model at an early stage, where the data feature resides in a raw energy distribution space rather than a box location feature space.
The result of the experiment shows that a reasonable fusion logic would significantly enhance the performance of the proposed RaLiBEV. While direct transformer fusion among feature maps can enhance performance, the implementation of interactive transformer-based concepts can yield superior results. This is largely due to the innovative querying fusion method, which introduces symmetry and unifying properties into the model.
To also highlight the benefits of multi-modality fusion, Table III presents a comparison of performance metrics between the LiDAR-only approach and the fusion of radar and LiDAR. It is evident from the comparison that neither RaLiBEV with LiDAR only nor PointPillars can match the performance levels achieved by the fusion of radar and LiDAR. This performance enhancement is particularly noticeable in foggy testing scenarios. When the model is trained on a “Clear+Foggy” dataset, RaLiBEV, in the radar and LiDAR fusion modality (denoted as “R+L”), exhibits exceptionally high performance, especially under the stringent requirements of a 0.8 IoU in “Foggy” test scenarios.
@@ IV-C3 Performance Comparison with State-of-the-Art Methods on Different Weather Conditions
Figure 6: Visualization of detection results of MVDNet and RaLiBEV.
Following MVDNet [24], model performance in different weather is tested in this paper as described in Table IV. The simulated foggy LiDAR data is generated by the code provided by MVDNet. This experiment is set as training using “Clear-only” and “Clear+Foggy” LiDAR and radar data, then testing under both “Clear” and “Foggy” data. Two peak performance models are chosen to prove their capability under this condition. One is RaLiBEV with GACHIPS, the other is RaLiBEV with DQMITBF. The performance of GACHIPS training under “Clear-only” data and testing under the “Foggy” scenario is better than DQMITBF’s. The performance of the horizontal comparison model after data enhancement is about 10%. GACHIPS increases from 84.6%, 82.4%, and 73.8% to 97.8%, 96.7%, and 93.7% under the “Foggy” test. This result exceeds the ST-MVDNet [25] over 20% at “Clear+Foggy” training and“Foggy” test scenario. Compared to the newest work Bi-LRFusion [53], RaLiBEV outperforms a large margin of 20.7% in IoU 0.8 requirement. Meanwhile, according to the results of the state-of-art work ST-MVDNet++ [52], RaLiBEV has made progress of 4.4%, 7.8% and 19.0% in IoU 0.5, 0.65 and 0.8 under “Clear+Foggy” training, “Foggy” test. This result represents that RaLiBEV is able to perform better accuracy in adverse weather such as fog. The fusion logic and label assignment strategy help the detector achieves higher precision.
Fig. 6 shows the ground-truth, the prediction from MVDNet in clear weather, the prediction from RaLiBEV in clear weather and foggy weather at the $3332^{th}$ and $4103^{th}$ frame of test dataset. The ground-truth bounding boxes are decorated with yellow color, and the predicted boxes by each model are shown with red color. All boxes use green lines to indicate the heading of the object. It is easy to find out that the prediction of MVDNet have 3 miss detection and 4 predicted boxes with reversed heading direction in the top row figure, and 1 false alarm and 2 reversed predicted boxes in the bottom row figure, though the weather is clear. In contrast, The RaLiBEV shows perfect detection results in clear weather, and only gives one false detection and no reversed heading prediction in both time frame under foggy weather.